18. A fair coin is flipped multiple times until it lands on heads. If the probability of landing on ( point) heads is 50%, what is the probability of first landing on heads on the fourth attempt? 00.625 0.500 00.412 00.382
Answer #1

P(head), p = 0.5

T stands for tail and H stands for head.

P(first heads on 4th attempt) = P(TTTH) = P(T) * P(T) * P(T) * P(H) = 0.5 * 0.5 * 0.5 * 0.5 = 0.0625